OpenStudy (anonymous):

f(x)= 4x-3 and g(x)= 1/x-9 Find the composite function f(g(x))

OpenStudy (anonymous):

f(g(x))= 4(1/x-9)-3

OpenStudy (anonymous):

I know that part, I'm not sure how to handle the fraction?

OpenStudy (anonymous):

ok (4/x-9)-3=0 4/x-9=3 3(x-9)=4 3x-27=4 3x=4+27 3x=31 x=31/3